XL Co.’s dividends are expected to grow at a 20% rate for the next 3 years, with the growth rate falling off to a constant 6% thereafter. If the required return is 14% and the company just paid a $3.10 dividend, what is the current price?XL Co.’s dividends are expected to grow at a 20% rate for the next 3 years, with the growth rate falling off to a constant 6% thereafter. If the required return is 14% and the company just paid a $3.10 dividend, what is the current price?
Answer #1

From the given information about XL Co.'s

required rate of return(Ke) = 14%,

For 1st 3 years

Growth rate = 20%

Dividend at time o = $3.1

So, D1 = 3.1*1.2 = $3.72

After 3 years

Constant growth rate = 6%

So, dividend for 1st 4 years will look like

Year 0 1 2 3 4
Dividend 3.1 3.72 4.46 5.36 5.68

1st calculating Horizon value of price at year 3 using Gordon Constant growth model

P3 = D4/(Ke-g) = 5.68/(0.14-0.06) = $70.98

Calculation price of share at year 0 using sum of PV of all dividends and PV of Horizon value of price at year 3

Price0 = D1/(1+Ke) + D2/(1+Ke)^2 + D3/(1+Ke)^3 + P3/(1+Ke)^3

Price0 = 3.72/1.14 + 4.46/1.14^2 + 5.36/1.14^3 + 70.98/1.14^3

Price0 = $58.22

current price of XL Co.'s stock = $58.22